OpenStudy (anonymous):

Adult tickets for a play cost $25 each and children's tickets cost $15 each. The total receipts were $7200 and the total attendance was 400. How many adults and how many children attended?

OpenStudy (jonnyvonny):

$25 for each adult and $15 for child, totaling a value we don't know. We know that adults and children came, the amount, unknown. \[25a+15c=y \left| \right| a+c=400\] Did they give us the total amount of money made?

OpenStudy (anonymous):

yes i just edited the question, sorry.

OpenStudy (jonnyvonny):

np

OpenStudy (jonnyvonny):

In which case:\[25a+15c=7,200\]\[a+c=400\] We can solve this using 1 or 2 ways; elimination or substitution, which do you want to do?

OpenStudy (anonymous):

I will try elimination, hold on...

OpenStudy (anonymous):

so the answer is: (120,280), or there were 120 adults and 280 children.

OpenStudy (jonnyvonny):

I got that as well. Good job :)...you know, of course, if we're both wrong then sheet >.>

OpenStudy (anonymous):

haha that would suck... should i write the answer the way I just wrote it or is there a better way to write it?

OpenStudy (jonnyvonny):

Write it out; don't use points. "280 children and 120 adults" works.

OpenStudy (anonymous):

okay, thank you!